The diagonal of a rectangle is a line drawn from one vertex to the opposite vertex. 5 mi and 10 mi
Find the length of the diagonal in the rectangle. Round to the nearest tenth.

Answer by ewatrrr(24785)


Hi
 Applying the Pythagorean Theorem to find the length of the diagonal
  Sides of the rectangle: 5 mi and 10 mi
    
      d^2 = 5mi^2 + 10mi^2 = 125
         d^2 = 125
           d = 11.2 mi   Round to the nearest tenth.


Important You are comfortable using your calculator
Wish You the Best in your Studies.